OpenStudy (calculusxy):

CAN SOMEONE HELP ME SOLVE THIS EQUATION GIVEN FROM A PREVIOUS POST?? {MEDAL!!!} (n+3)(x-1)=216

OpenStudy (agent47):

solve for?

OpenStudy (anonymous):

n or x ?

OpenStudy (calculusxy):

n

OpenStudy (agent47):

(n+3)=216/(x-1) n = 216/(x-1) - 3

OpenStudy (calculusxy):

Then what can I do?

OpenStudy (agent47):

\[n = \frac{216-3(x-1)}{x-1}\]

OpenStudy (agent47):

\[n=\frac{216-3x+3}{x-1}=\frac{219-3x}{x-1}\]

OpenStudy (agent47):

\[n=\frac{3(73-x)}{x-1}\]

OpenStudy (agent47):

That's about it..

OpenStudy (calculusxy):

So what does n equal to?

OpenStudy (radar):

Depends on value of x, if x = 0, then n=-219, or if x =1, n is undefined. etc.

OpenStudy (calculusxy):

Well I had this one question that I needed help on. Someone did, but then they gave me an equation which I had no idea on how to solve it. But can you please help me and explain all this material to me?

OpenStudy (calculusxy):

Ella spent her whole allowance of $2.00, plus the 16¢ she had left over from last week, on bubble gum. If the pieces of gum had been a penny cheaper, she would have received three more pieces than she did. How many pieces did she actually buy?

OpenStudy (radar):

Use a little dab of algebra for this bubble gum problem. Given cost is 216 cents for total gum purchased Let x = the number of pieces of gum at original price. then 216/x = price of gum at original price. Now for the if portion. if price per gum is 1 cent less then would of cost (216/x)-1 per piece (give) the if or new quantity purchased would be x + 3 (given) From this we can state (mathematically): (x+) [(216/x)-1)] = 216 now solve listed are some intermediate steps \[x ^{2}+3x-648 =0\] (x+27)(x-24) = 0 x=24 getting the gum a 9 cents a piece. if x = 27, if gum sold at 8 cents a piece. the Lesson being don't spend your entire allowance on bubble gum.